OpenStudy (anonymous):

For questions 1-2, convert to logarithmic equations. For example, the logarithmic form of "23 = 8" is "log2 8 = 3". 1. 9 ^3/2 = 27 2. e^x = 7

OpenStudy (anonymous):

log 9-27 = 3/2

OpenStudy (anonymous):

i need steps? so i can figure how to do the rest?

OpenStudy (anonymous):

its rearanging the exponents to log, for example, 2^2=4 same as log (base of the number (2) THEN- THE ANSWER of 2^2= (4), and the power is behind the equal sign, SO log2-4=2 same as (2^2=4)

OpenStudy (anonymous):

yea, i don't understand that

OpenStudy (anonymous):

ok, ill give you some examples

OpenStudy (anonymous):

log 5-25=2 5^2=25 log 8-64=2 8^2=64 log x-y=e x^e=y you arent changing any numbers, just rearanging them